Circular No. 7030

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998ed IN ANONYMOUS GALAXY J. Mueller reports her discovery of an apparent supernova (mag ∼ 18)

on a plate taken on Sept. 16 with the 1.2-m Oschin Schmidt Telescope by K. Rykoski in the course of the Palomar Outer Solar System Ecliptic Survey. SN 1998ed is located at α = 23^{h}00^{m}14^{s}.64, δ = −13^{o}13′58′′.2 (equinox 2000.0), which is 6′′ west and 2′′.7 south of the center of the host galaxy. The new object is also visible on plates taken on Sept. 14 and 15 UT. There is no object at this position on the Digital Sky Survey.

SUPERNOVA 1998ee IN ANONYMOUS GALAXY J. Maza, Department of Astronomy, University of Chile, reports the

discovery by Marina Wischnjewsky, on a T-Max 400 film taken by L. E. González using the Maksutov telescope at Cerro El Roble on Oct. 14.255 UT, of a supernova (B ∼ 17.5) located at α = 1^{h}53^{m}31^{s}.33, δ = −53^{o}58′19′′.5 (equinox 2000.0), which is 5′′.1 east and 1′′.9 north of the nucleus of the host spiral galaxy. The object was confirmed via B and V direct images taken by D. González on Oct. 16.46 using the YALO telescope at Cerro Tololo. SN 1998ee is not visible on films obtained with the Cerro El Roble telescope on Sept. 27.245 through clouds (B > 17) and on Sept. 15.256 (B > 19).

PSR 1937+21 M. Takahashi and S. Shibata, Yamagata University; K. Torii, National

Space Development Agency of Japan; Y. Saito, Institute of Space and Astronautical Science; and N. Kawai, Institute of Physical and Chemical Research, report: “The x-ray counterpart to the fastest millisecond pulsar PSR 1937+21 was detected with ASCA on 1997 Dec. 16.0–17.1 UT, setting a strong constraint on the emission model for the pulsar radiation. The x-ray source pulsed at the same period as found in the radio bands, and had a hard spectrum consistent with a power law with photon index 0.9 and flux 5.9 × 10^{−13} erg cm^{−2} s^{−1} (2–10 keV). Most of the x-ray flux (50–90 percent in the band 1.7-6.5 keV) was contained in the narrow primary peak with width probably < 130 μs, and with an indication of an interpulse.”

COMET C/1998 M5 (LINEAR) Visual m_{1} estimates by R. Bouma, Groningen, The Netherlands (0.25-

m refl.): Sept. 22.85 UT, 10.9; 23.87, 11.0; Oct. 12.89, 10.8; 14.83, 10.8.

1998 October 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green